# Multimodal Data Integration for Oncology in the Era of Deep Neural Networks: A Review

### v2

@@ 2.3 Data Fusion Stages
The intermediate fusion approach involves training separate models on each modality and then combining the outputs of these models to make a final prediction [16]. This approach is suitable when the data modalities are independent of each other and cannot be easily combined at the feature level. The main challenge with intermediate fusion is selecting an appropriate method for combining the output of the separate models. This can be done by averaging the separate models’ output or using a weighted sum.
@@
In late fusion, the output of each model is used to make a separate decision, which is then combined to make a final decision. This approach is suitable when the modalities provide complementary information but are not necessarily independent of each other. The main challenge with late fusion is selecting an appropriate method for combining individual predictions. This can be done using majority voting, weighted voting, or employing other ML models.
2.4 Neural Network Architectures on Oncology Data
@@
ML and DL techniques are pervasive and effective in oncology research [39]. Cahall et al. [40] used Inception modules on U-Net for brain tumor segmentation. Syed et al. [41] used a Random Forest classifier to fuse radiology image representations learned from the singular value decomposition method with the textual annotation representation learned from the fastText algorithm for prostate and lung cancer patients. Liu et al. [42] proposed a hybrid DL framework for combining breast cancer patients’ genomic and pathology data using fully-connected (FC) network for genomic data, CNN for radiology data, and Simulated Annealing algorithm for late fusion. Multiview multimodal network (MVMM-Net) [43] combined two different modalities (low-energy and dual-energy subtracted) from contrast-enhanced spectral mammography images, each learned through CNN and late-fusion through FC network in breast cancer detection task. Yap et al. [44] used a late-fusion method to fuse image representations from ResNet50 and clinical representations from random forest for a multimodal skin lesion classification task. An award-winning work [45] on brain tumor grade classification followed a late-fusion (concatenation) from CNNs on radiology and pathology images separately and classified using logistic regression algorithm.
@@
The single-cell unimodal data alignment is one technique in MML. Jansen et al. devised an approach (SOMatic) to combine ATAC-seq regions with RNA-seq genes using self-organizing maps [46]. Single-Cell data Integration via Matching (SCIM) [47] matched cells in multiple datasets in low-dimensional latent space using autoencoder. Graph-linked unified embedding (GLUE) [48] model learned regulatory interactions across omics layers and aligned the cells using variational autoencoders. These aforementioned methods cannot incorporate high-order interactions among cells or different modalities. Single-cell data integration using multiple modalities is mostly based on auto-encoders (scDART [49], Cross-modal Autoencoders [50], Mutual Information Learning for Integration of Single Cell Omics Data (SMILE) [51]).
3 GNNs in Multimodal Learning
@@
Graphs are commonly used to represent the relational connectivity of any system that has interacting entities[52]. Graphs have been used in various fields, such as to study brain networks [53], driving maps [54], and the structure of DNNs themselves [55]. Graph neural networks (GNNs) are a type of neural network specifically designed to process data represented as a graph [56], which makes them well-suited for analyzing multimodal oncology data [12]. In this section, we first introduce the graph data with its types and tasks, then we discuss the various graph learning algorithms, followed by a detailed review of the graph learning methods used on unimodal and multimodal oncology data tasks.
@@ 3.1 The Graph Data
[…]
@@ 3.1.1 Graph Types
As illustrated in Fig. 5(a), the common types of graphs include undirected, directed, homogeneous, heterogeneous, static, dynamic, unattributed, and attributed. Undirected graphs comprise undirected edges, that is, the direction of relation is not important between any ordered pair of nodes. In the directed graphs, the nodes have a directional relationship(s). Homogeneous graphs have the same type of nodes, whereas heterogeneous graphs have different types of nodes within a single graph [57]. Static graphs do not change over time with respect to the existence of edges and nodes, whereas dynamic graphs change over time resulting in changes to structure, attributes, and node relationships. Unattributed graphs have unweighted edges indicating that the weighted value for all edges in a graph is the same, i.e., $1$ if present, 00 if absent. Attributed graphs have different edge weights that capture the strength of relational importance [56].
Figure 5: (a) The commonly occurring graph types: undirected/ directed, homogeneous/ heterogeneous, dynamic/ static, attributed (edges)/ Unattributed. (b) Three tasks on graph data: node-level, link-level, and graph-level tasks. (c) Various categories of representation learning for graphs are presented.
3.1.2 Tasks on Graphs
In Fig. 5(b), we present three major types of tasks defined on graphs, including:
  • Node-level tasks. These may include node classification, regression, clustering, attributions, and generation.
  • Edge-level task. Edge classification and prediction (presence or absence) are two common edge-level tasks.
  • Graph-level tasks. These tasks involve predictions on the graph level, such as graph classification and generation.
3.2 Learning on Graphs
Representing data as graphs can enable capturing and encoding the relationships among entities of the samples [58]. Based on the way the nodes are encoded, the data representation learning on graphs can be categorized into the traditional (or shallow) and deep neural network-based methods, as illustrated in Fig. 5(c). The reader is referred to [59, 58] for details on these methods.
@@ 3.2.1 Traditional (Shallow) Methods
3.2.2 Deep Methods - GNNs
GNNs are gaining popularity in the ML community, as evident from Fig. 1. GNNs are DL techniques defined for graphs where information aggregation from the neighborhood is fused into a node’s representation. Traditional DL methods such as CNNs and their variants have shown remarkable success in processing the data in Euclidean space; however, they fail to perform well when faced with non-Euclidean or relational datasets. Compared to CNNs, where the locality of the nodes in the input is fixed, GNNs have no canonical ordering of the neighborhood of a node and can learn the given task for any permutation of the input data, as depicted in Fig. 6. GNNs often employ a message-passing mechanism in which a node’s representation is derived from its neighbors’ representations via a recursive computation. The message passing for a GNN is given as follows:
@@ 3.2.2 Deep Methods - GNNs
Figure 6: Image vs. graph convolutions. The canonical order of the input is important in CNNs, whereas in GNNs, the order of the input nodes is not important. This essentially makes CNNs a subset of GNNs [60].
where $h_{v}^{(l+1)}$ is the updated embedding of node $v$ after $l$+$1$ layer, $\sigma$ is the non-linear function (such as rectified linear unit or ReLU), $h_{u}^{(l)}$ and $h_{v}^{(l)}$ represent the embeddings of nodes $u$ and $v$ at layer $l$. $W_{l}$ and $B_{l}$ are the trainable weight matrices for neighborhood aggregation and (self)hidden vector transformation, respectively. The message passing can encode high-order structural information in node embedding through multiple aggregation layers. GNNs smooth the features by aggregating neighbors’ embedding and filter eigenvalues of graph Laplacian, which provides an extra denoising mechanism [61]. GNNs comprise multiple permutation equivariant and invariant functions, and they can handle heterogeneous data [62]. As described earlier, traditional ML models deal with Euclidean data. In oncology data, the correlations may not exist in Euclidean space; instead, its features may be highly correlated in the non-Euclidean space [63]. Based on the information fusion and aggregation methodology, GNNs-based deep methods are classified into the following:
  • Recurrent GNNs: RecGNNs are built on top of the standard Recurrent Neural Network (RNN) by combining with GNN. RecGNNs can operate on graphs having variable sizes and topologies. The recurrent component of the RecGNN captures temporal dependencies and learns latent states over time, whereas the GNN component captures the local structure. The information fusion process is repeated a fixed number of times until an equilibrium or the desired state is achieved [64]. RecGNNs employ the model given by:
    $\mathbf{h}_{v}^{(l+1)}=\text{RecNN}\left(\mathbf{h}_{u}^{(l)},\mathbf{Msg}_{N(v)}^{(l)}\right)$ (2)
@@
  • Convolutional GNNs: ConvGNNs undertake the convolution operation on graphs by aggregating neighboring nodes’ embeddings through a stack of multiple layers. ConvGNNs use the symmetric and normalized summation of the neighborhood and self-loops for updating the node embeddings given by:
    $\mathbf{h}_{v}^{(l+1)}=\sigma\left(W_{l}\sum_{u\in N(v)\cup v}\frac{\mathbf{h}_{v}}{\sqrt{|N(v)||N(u)|}}\right)$ (3)
    The ConvGNN can be spatial or spectral, depending on the type of convolution they implement. Convolution in spatial ConvGNNs involves taking a weighted average of the neighboring vertices. Examples of spatial ConvGNNs include GraphSAGE[64], Message Passing Neural Network (MPNN)[65], and Graph Attention Network (GAT) [66]. The spectral ConvGNNs operate in the spectral domain by using the eigendecomposition of the graph Laplacian matrix. The convolution operation is performed on the eigenvalues, which can be high-dimensional. Popular spectral ConvGNNs are ChebNet [67] and Graph Convolutional Network (GCN)[68]. An interesting aspect of these approaches is representational containment, which is defined as:
    $convolution\subseteq attention\subseteq message\_passing.$
  • Graph Auto-Encoder Networks: GAEs are unsupervised graph learning networks for dimensionality reduction, anomaly detection, and graph generation. They are built on top of the standard auto-encoders to work with graph data. The encoder component of the GAE maps the input graph to a low-dimensional latent space, while the decoder component maps the latent space back to the original graph [69].
  • Graph Adversarial Networks (GraphANs): Based on Generative Adversarial Networks, GraphANs are designed to work with graph-structured data and can learn to generate new graphs with similar properties to the input data. The generator component of the GraphAN maps a random noise vector to a new graph, while the discriminator component tries to distinguish between the generated vs. the actual input. The generator generates graphs to fool the discriminator, while the discriminator tries to classify the given graph as real or generated.
  • Other GNNs: Other categories of GNNs may include scalable GNNs [70], dynamic GNNs [71], hypergraph GNNs [72], heterogeneous GNNs [73], and many others [74].
@@
3.3 Unimodal Oncology Learning using GNNs
Here, we explore how GNNs have been used to analyze single-modality oncology data. Single modality refers to using one resolution or type of cancer data, such as pathology, radiology, or molecular. A detailed survey on the use of GNNs for both oncology and non-oncology data is presented in [77].
3.3.1 Pathology
@@
Traditionally, CNN-based models are predominantly used to learn features from digital pathology data. However, unlike GNNs, CNNs fail to capture the global contextual information which is important in the tissue phenotypical and structural micro and macro environment [78]. For using histology images in GNNs, the cells, tissue regions, or image patches are depicted as nodes. The relations and interactions among these nodes are represented as (un)weighted edges. Usually, a graph of the patient histology slide is used along with a patient-level label for training a GNN, as illustrated in Fig. 7(a). Here we review a few GNNs-based pathology publications representative of a trove of works in this field. Histographs [79] used breast cancer histology data to distinguish cancerous and non-cancerous images. Pre-trained VGG-UNet was used for nuclei detection, micro-features of the nuclei were used as node features, and Euclidean distance among nuclei was incorporated as edge features. The resulting cell-graphs were used to train the GCN-based robust spatial filtering (RSF) model and showed superior performance compared to the CNN-based classification. Wang et al. [80] analyzed grade classification in tissue micro-arrays of prostate cancer using the weakly-supervised technique on a variant of GraphSAGE with self-attention pooling (SAGPool). Cell-Graph Signature ($CG_{signature}$[81] predicted patient survival in gastric cancer using cell-graphs of multiplexed immunohistochemistry images processed through two types of GNNs (GCNs and GINs) with two types of pooling (SAGPool, TopKPool). Besides above mentioned cell-graphs, there is an elaborate review of GNN-based tissue-graphs or patch-graphs methods implemented on unimodal pathology cancer data given in [78]. Instead of individual cell- and tissue-graphs, a combination of the multilevel information in histology slides can help understand the intrinsic features of the disease.
3.3.2 Radiology
@@
GNNs have been used in radiology-based cancer data for segmentation, classification, and prediction tasks, especially on X-rays, mammograms, MRI, PET, and CT scans. Fig. 7(b) illustrates a general pipeline of using radiology-based data to train GNNs. Here we give a non-exhaustive review of GNNs-based works on radiological oncology data as a single modality input. Mo et al. [82] proposed a framework that improved the liver cancer lesion segmentation in the MRI-T1WI scans through guided learning of MRI-T2WI modality priors. Learned embeddings from fully convolutional networks on separate MRI modalities are projected into the graph domain for learning by GCNs through the co-attention mechanism and finally to get the refined segmentation by re-projection. Radiologists usually review radiology images by zooming into the region of interest (ROIs) on high-resolution monitors. Du et al. [83] used a hierarchical GNN framework to automatically zoom into the abnormal lesion region of the mammograms and classify breast cancer. The pre-trained CNN model extracts image features, whereas a GAT model is used to classify the nodes for deciding whether to zoom in or not based on whether it is benign or malignant. Based on the established knowledge that lymph nodes (LNs) have connected lymphatic system and LNs cancer cells spread on certain pathways, Chao et al. [84] proposed a lymph node gross tumor volume learning framework. The framework was able to delineate the LN appearance as well as the inter-LN relationship. The end-to-end learning framework was superior to the state-of-the-art on esophageal cancer radiotherapy dataset. Tian et al. [85] suggested interactive segmentation of MRI scans of prostate cancer patients through a combination of CNN and two GCNs. CNN model outputs a segmentation feature map of MRI, and the GCNs predict the prostate contour from this feature map. Saueressig et al. [86] used GNNs to segment brain tumors in 3D MRI images, formed by stacking different modalities of MRI (T1, T2, T1-CE, FLAIR) and representing them as supervoxel graph. The authors reported that GraphSAGE-pool was best for segmenting brain tumors. Besides radiology, a parallel field of radiomics has recently gained attraction. Radiomics is the automated extraction of quantitative features from radiology scans. A survey of radiomics and radiogenomic analysis on brain tumors is presented in[76].
3.3.3 Molecular Data

### v3

@@ 1.3.2 Intermediate Fusion
Intermediate fusion involves training separate models for each data modality and then combining the outputs of these models for inference/prediction [34]. This approach is suitable when the data modalities are independent of each other and cannot be easily combined at the feature level using average, weighted average, or other methods. The main challenge with intermediate fusion is selecting an appropriate method for combining the output of different models.
1.3.3 Late Fusion
@@
In late fusion, the output of each modality-specific model is used to make a decision independently. All decisions are later combined to make a final decision. This approach is suitable when the modalities provide complementary information but are not necessarily independent of each other. The main challenge with late fusion is selecting an appropriate method for combining individual predictions. This can be done using majority voting, weighted voting, or employing other ML models.
1.4 MML for Oncology Datasets
@@
Syed et al. [56] used a Random Forest classifier to fuse radiology image representations learned from the singular value decomposition method with the textual annotation representation learned from the fastText algorithm for prostate and lung cancer patients. Liu et al. [57] proposed a hybrid DL framework for combining breast cancer patients’ genomic and pathology data using fully-connected (FC) network for genomic data, CNN for radiology data and a Simulated Annealing algorithm for late fusion. Multiview multimodal network (MVMM-Net) [58] combined 2 different modalities (low-energy and dual-energy subtracted) from contrast-enhanced spectral mammography images, each learned through CNN and late-fusion through FC network in breast cancer detection task. Yap et al. [59] used a late-fusion method to fuse image representations from ResNet50 and clinical representations from a random forest model for a multimodal skin lesion classification task. An award-winning work [60] on brain tumor grade classification adopted the late-fusion method (concatenation) for fusing outputs from two CNNs (radiology and pathology images).
@@
The single-cell unimodal data alignment is one technique in MML. Jansen et al. devised an approach (SOMatic) to combine ATAC-seq regions with RNA-seq genes using self-organizing maps [61]. Single-Cell data Integration via Matching (SCIM) [62] matched cells in multiple datasets in low-dimensional latent space using autoencoder (AEs). Graph-linked unified embedding (GLUE) [63] model learned regulatory interactions across omics layers and aligned the cells using variational AEs. These aforementioned methods cannot incorporate high-order interactions among cells or different modalities. Single-cell data integration using multiple modalities is mostly based on AEs (scDART [64], Cross-modal Autoencoders [65], Mutual Information Learning for Integration of Single Cell Omics Data (SMILE) [66]).
2 Graph Neural Networks (GNNs) in Multimodal Learning
@@
Graphs are commonly used to represent the relational connectivity of any system that has interacting entities [67]. Graphs have been used in various fields, such as to study brain networks [68], analyze driving maps [69], and explore the structure of DNNs themselves [70]. GNNs are specifically designed to process data represented as a graph [71], which makes them well-suited for analyzing multimodal oncology data as each data modality (or sub-modality) can be considered as a single node and the structures/patterns that exist between data modalities can be modeled as edges [32].
@@ 2.1 The Graph Data
A graph is represented as $G$=$(V,E)$ having node-set $V$=$\{v_{1},v_{2},...,v_{n}\}$, where node $v$ has feature vector xv, and edge set $E$=$\{(v_{i},v_{j})\mid v_{i},v_{j}\in V\}$. The neighborhood of node $v$ is defined as $N(v)$=$\{u\mid(u,v)\in E\}$.
@@ 2.1.1 Graph Types
As illustrated in figure 4(a), the common types of graphs include undirected, directed, homogeneous, heterogeneous, static, dynamic, unattributed, and attributed.Undirected graphs comprise undirected edges, i.e., the direction of relation is not important between any ordered pair of nodes. In the directed graphs, the nodes have a directional relationship(s). Homogeneous graphs have the same type of nodes, whereas heterogeneous graphs have different types of nodes within a single graph [72]. Static graphs do not change over time with respect to the existence of edges and nodes. In contrast, dynamic graphs change over time, resulting in changes in structure, attributes, and node relationships. Unattributed graphs have unweighted edges, indicating that the weighted value for all edges in a graph is the same, i.e., $1$ if present, 00 if absent. Attributed graphs have different edge weights that capture the strength of relational importance [71].
Figure 4: (a) The commonly occurring graph types are presented, including (1) undirected and directed, (2) homogeneous and heterogeneous, (3) dynamic and static, (4) attributed (edges) and unattributed. (b) Three different types of tasks performed using the graph data are presented and include (1) node-level, (2) link-level, and (3) graph-level analyses. (c) Various categories of representation learning for graphs are presented.
2.1.2 Tasks for Graph Data
In figure 4(b), we present 3 major types of tasks defined on graphs, including (1) node-level tasks - these may include node classification, regression, clustering, attributions, and generation, (2) edge-level task - edge classification and prediction (presence or absence) are 2 common edge-level tasks, (3) graph-level tasks - these tasks involve predictions on the graph level, such as graph classification and generation.
2.2 ML for Graph Data
Representing data as graphs can enable capturing and encoding the relationships among entities of the samples [73]. Based on the way the nodes are encoded, representation learning on graphs can be categorized into the traditional (or shallow) and DNN-based methods, as illustrated in Figure 4(c)  [74, 73].
@@ 2.2.1 Traditional (Shallow) Methods
2.2.2 DNN-based Methods - GNNs
GNNs are gaining popularity in the ML community, as evident from figure 1. In GNNs, the information aggregation from the neighborhood is fused into a node’s representation. Traditional DL methods such as CNNs and their variants have shown remarkable success in processing the data in Euclidean space; however, they fail to perform well when faced with non-Euclidean or relational datasets. Compared to CNNs, where the locality of the nodes in the input is fixed, GNNs have no canonical ordering of the neighborhood of a node. They can learn the given task for any permutation of the input data, as depicted in figure 5. GNNs often employ a message-passing mechanism in which a node’s representation is derived from its neighbors’ representations via a recursive computation. The message passing for a GNN is given as follows:
@@ 2.2.2 DNN-based Methods - GNNs
Figure 5: Convolution operation for graphs vs. image data. The canonical order of the input is important in CNNs, whereas in GNNs, the order of the input nodes is not important. From the convolution operation perspective, CNNs can be considered a subset of GNNs [75].
where $h_{v}^{(l+1)}$ is the updated embedding of node $v$ after $l$+$1$ layer, $\sigma$ is the non-linear function (e.g., rectified linear unit or ReLU), $h_{u}^{(l)}$ and $h_{v}^{(l)}$ represent the embeddings of nodes $u$ and $v$ at layer $l$. $W_{l}$ and $B_{l}$ are the trainable weight matrices for neighborhood aggregation and (self)hidden vector transformation, respectively. The message passing can encode high-order structural information in node embedding through multiple aggregation layers. GNNs smooth the features by aggregating neighbors’ embedding and filter eigenvalues of graph Laplacian, which provides an extra denoising mechanism [76]. GNNs comprise multiple permutation equivariant and invariant functions, and they can handle heterogeneous data [77]. As described earlier, traditional ML models deal with Euclidean data. In oncology data, the correlations may not exist in Euclidean space; instead, its features may be highly correlated in the non-Euclidean space [78]. Based on the information fusion and aggregation methodology, GNNs-based deep methods are classified into the following:
Recurrent GNNs
RecGNNs are built on top of the standard Recurrent Neural Network (RNN) by combining with GNN. RecGNNs can operate on graphs with variable sizes and topologies. The recurrent component of the RecGNN captures temporal dependencies and learns latent states over time, whereas the GNN component captures the local structure. The information fusion process is repeated a fixed number of times until an equilibrium or the desired state is achieved [79]. RecGNNs employ the model given by:
$\mathbf{h}_{v}^{(l+1)}=\text{RecNN}\left(\mathbf{h}_{u}^{(l)},\mathbf{Msg}_{N(v)}^{(l)}\right),$ (2)
@@
Convolutional GNNs
ConvGNNs undertake the convolution operation on graphs by aggregating neighboring nodes’ embeddings through a stack of multiple layers. ConvGNNs use the symmetric and normalized summation of the neighborhood and self-loops for updating the node embeddings given by:
$\mathbf{h}_{v}^{(l+1)}=\sigma\left(W_{l}\sum_{u\in N(v)\cup v}\frac{\mathbf{h}_{v}}{\sqrt{|N(v)||N(u)|}}\right).$ (3)
[…]
Graph Auto-Encoder Networks (GAEs)
GAEs are unsupervised graph learning networks for dimensionality reduction, anomaly detection, and graph generation. They are built on top of the standard AEs to work with graph data. The encoder component of the GAE maps the input graph to a low-dimensional latent space, while the decoder component maps the latent space back to the original graph [84].
Graph Adversarial Networks (GraphANs)
Based on Generative Adversarial Networks, GraphANs are designed to work with graph-structured data and can learn to generate new graphs with similar properties to the input data. The generator component of the GraphAN maps a random noise vector to a new graph, while the discriminator component tries to distinguish between the generated vs. the actual input. The generator generates graphs to fool the discriminator, while the discriminator tries to classify the given graph as real or generated.
Other GNNs
Other categories of GNNs may include scalable GNNs [85], dynamic GNNs [86], hypergraph GNNs [87], heterogeneous GNNs [88], and many others [89].
@@
2.3 GNNs and ML using Unimodal Oncology Datasets
2.3.1 Pathology Datasets
@@
Traditionally, CNN-based models are used to learn features from digital pathology data. However, unlike GNNs, CNNs fail to capture the global contextual information important in the tissue phenotypical and structural micro and macro environment [92]. For using histology images in GNNs, the cells, tissue regions, or image patches are depicted as nodes. The relations and interactions among these nodes are represented as (un)weighted edges. Usually, a graph of the patient histology slide is used along with a patient-level label for training a GNN, as illustrated in Figure 6(a). Here, we review a few GNN-based pathology publications representative of a trove of works in this field. Histographs [93] used breast cancer histology data to distinguish cancerous and non-cancerous images. Pre-trained VGG-UNet was used for nuclei detection, micro-features of the nuclei were used as node features, and Euclidean distance among nuclei was incorporated as edge features. The resulting cell graphs were used to train the GCN-based robust spatial filtering (RSF) model, which performed superior to the CNN-based classification. Wang et al. [94] analyzed grade classification in tissue micro-arrays of prostate cancer using the weakly-supervised technique on a variant of GraphSAGE with self-attention pooling (SAGPool). Cell-Graph Signature ($CG_{signature}$[95] predicted patient survival in gastric cancer using cell-graphs of multiplexed immunohistochemistry images processed through two types of GNNs (GCNs and GINs) with two types of pooling (SAGPool, TopKPool). Besides the above-mentioned cell graphs, there is an elaborate review of GNN-based tissue graphs or patch-graphs methods implemented on unimodal pathology cancer data given in [92]. Instead of individual cell- and tissue-graphs, a combination of the multilevel information in histology slides can help understand the intrinsic features of the disease.
2.3.2 Radiology Datasets
@@
GNNs have been used in radiology-based cancer data for segmentation, classification, and prediction tasks, especially on X-rays, mammograms, MRI, PET, and CT scans. Figure 6(b) illustrates a general pipeline of using radiology-based data to train GNNs. Here we give a non-exhaustive review of GNNs-based works on radiological oncology data as a single modality input. Mo et al. [96] proposed a framework that improved the liver cancer lesion segmentation in the MRI-T1WI scans through guided learning of MRI-T2WI modality priors. Learned embeddings from fully convolutional networks on separate MRI modalities are projected into the graph domain for learning by GCNs through the co-attention mechanism and finally to get the refined segmentation by re-projection. Radiologists usually review radiology images by zooming into the region of interest (ROIs) on high-resolution monitors. Du et al. [97] used a hierarchical GNN framework to automatically zoom into the abnormal lesion region of the mammograms and classify breast cancer. The pre-trained CNN model extracts image features, whereas a GAT model is used to classify the nodes for deciding whether to zoom in or not based on whether it is benign or malignant. Based on the established knowledge that lymph nodes (LNs) have connected lymphatic system and LNs cancer cells spread on certain pathways, Chao et al. [98] proposed a lymph node gross tumor volume learning framework. The framework was able to delineate the LN appearance as well as the inter-LN relationship. The end-to-end learning framework was superior to the state-of-the-art on esophageal cancer radiotherapy dataset. Tian et al. [99] suggested interactive segmentation of MRI scans of prostate cancer patients through a combination of CNN and two GCNs. CNN model outputs a segmentation feature map of MRI, and the GCNs predict the prostate contour from this feature map. Saueressig et al. [100] used GNNs to segment brain tumors in 3D MRI images, formed by stacking different modalities of MRI (T1, T2, T1-CE, FLAIR) and representing them as supervoxel graph. The authors reported that GraphSAGE-pool was best for segmenting brain tumors. Besides radiology, a parallel field of radiomics has recently gained attraction. Radiomics is the automated extraction of quantitative features from radiology scans. A survey of radiomics and radiogenomic analysis on brain tumors is presented in[91].
2.3.3 Molecular Datasets